### v1

@@ 3.3 Image Retrieval by Overlap Scores
The retrieval pipeline is visualized in Fig. 4. In the online phase, we obtain the patch and, optionally, the global embeddings of the query image. Next, for each query image patch, we perform a radius search in the database of patch embeddings to determine all potentially overlapping candidates. Setting a radius threshold $\epsilon$ is easy as the employed cosine similarity falls between 0 and 1. We tune this threshold during training and use a fixed one in all our experiments. Note that k-nearest-neighbors algorithm is insufficient in our case as we would like to find all potentially overlapping patches to determine image overlaps. Such a radius search can be efficiently performed by standard bounding structures, e.g., kd-tree, and other implementation tricks. Optionally, we also use the global descriptor to prefilter the images to a shortlist.
To obtain the top-$k$ images, we perform voting such that the overlap calculation becomes
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\llbracket sim(p,q)>\epsilon\rrbracket,$ (8)
@@
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\max(sim(p,q)-\epsilon,0).$ (9)
Weighted Voting. To highlight the distinguishable query patches, we propose to weight the query patches from each image using a bag of visual words approach [16]. Specifically, we apply TF-IDF [58] weights to all the query patches, assigning lower weights to patches that frequently appear in queries but are rare in the matched database patches. As illustrated below, it is preferable to have more neighbor patches if fewer images share the same neighborhood with the query patch. For each patch, the weight $t_{i}$ is calculated as follows:
@@
where $N$ represents the total number of database images, $n_{d}$ is the total number of patches in each image, 256 in our case, with a patch size of (14, 14) on the resized image size of 224. Parameter $n_{id}$ is the number of neighbor patches in database images, defined as the cosine similarities of the patch descriptors within a specified radius, and $n_{i}$ denotes the number of database images that contain at least one neighbor of the query patch. Patches that frequently appear are deemed less important and thus are down-weighted. Weight $t_{i}$ for each query patch is incorporated into the overlap score computation as follows: $t_{i}*O(I_{i},I_{j})$. Note that the map weights are precomputed during the mapping phase.
@@ 4 Experimental Results
Our experiments concentrate on estimating two-view epipolar geometry for pairs of retrieved images. The model is trained on 153 scenes from the MegaDepth dataset [39], which includes Internet-sourced images and depths reconstructed via Structure from Motion (SfM) and multi-view stereo techniques. Beyond the MegaDepth testing scenes, we evaluate its generalization on 11 scenes from PhotoTourism [71], the 13 scenes from ETH3D [67], and on the InLoc dataset [74]. We benchmark the proposed method against several competitors, including the foundational DINO V2, CNN-based retrieval methods like NetVLAD [2] and CosPlace [5], as well as the state-of-the-art (SOTA) universal retrieval method AnyLoc [33] utilizing the same DINO V2 backbone as we do. Besides the global-only retrieval methods, we compare VOP with the SOTA reranking-based methods: ViT-based $R^{2}$ormer [90] and CNN-based Patch-NetVLAD [28]. Also, we show the results of CosPlace when fine-tuned on the same dataset we train on.
@@
Training Details. The models underwent training on the 153 MegaDepth scenes, employing contrastive loss to differentiate between negative and positive patch pairs. Each scene contributes 150 positive and 150 negative image pairs to the training set, standardized to an image resolution of $224\times 224$ pixels. We use the DINO V2 pre-trained model with a patch size of $14$ and extract $256$-dimensional descriptors. Negative images, characterized by zero overlaps (e.g., sourced from distinct scenes), contrast with positive pairs, which are stratified into three groups based on overlap scores ranging from 0.1 to 0.7 for random selection. Notably, the contrastive loss framework is predicated on patch-level negative/positive distinctions rather than entire images. Following the methodology of LightGlue [41], patch-level labels are generated from depth information and visual features, utilizing the GlueFactory library [52] for implementation.
@@
The training was conducted over 30 epochs with a batch size of 64, employing matching recalls as the primary criterion for best checkpoint selection and utilizing a learning rate of $41e-$ 4. The voting scheme and the radius for neighbor searching were optimized on the validation set, which included a single scene for this purpose. During inference, scenes “0015” (comprising approximately 0.4K images) and “0022” (encompassing approximately 0.5K images) from MegaDepth were designated for testing, enabling a focused evaluation of the model performance on these specific subsets.
@@
Moreover, we report the average number of inliers across image pairs as detected by the retrieval methods. It is important to note that in scenarios where $k>1$, for each query, we follow the practical approach and select only the image pair that has the highest number of inliers. This comprehensive evaluation allows us to thoroughly assess the performance not just in retrieving relevant images but also in accurately estimating their relative poses.
@@ 4.1 Pose Estimation on Megadepth
Table 1 presents the results on the MegaDepth dataset, highlighting the performance of two variants of our method. For VOP, we select the model that maximizes the pose AUC on the validation set. For VOP-recall, we choose the model that maximizes retrieval recall. This experiment highlights that these are two different objectives, and “recall” should not be used for retrieval if the final goal is localization or reconstruction. Retrieval recall provides the most similar images, but they may not be suitable for pose estimation due to short baselines. The proposed method secures the highest scores for the @10 criterion and exhibits the lowest median pose errors, thereby suggesting that recall alone may not be a comprehensive metric for evaluating the efficacy of VPR methodologies. Notably, the VOP-recall variant demonstrates superior accuracy compared to other methods when tailored to enhance recall, which underscores the limitations of recall as a sole metric.
The underlined results mark the optimal configuration for relative pose estimation accuracy, evidencing the superior performance of the proposed VOP approach when utilizing the top-5 images. This configuration outperforms the global-only retrieval baselines (the first four rows) on this dataset, reinforcing the significance of selecting appropriate metrics and model configurations tailored to the specific demands of real-world downstream applications. Besides, rows 5 and 6 in Table 1 show our method outperforms the recent reranking approaches $R^{2}$ormer [90] and Patch-NetVLAD [28] on Megadepth. Both lead to comparable results in top-10 retrieval but worse in the others.
Recall (%) $\uparrow$ AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10 1 5 10
AnyLoc [33] 16.7 24.4 29.5 60.8 65.3 64.1 2.38 2.26 2.35 220.5 273 287.5
CosPlace [5] 19.3 31.2 37.5 54.5 53.5 64.3 3.14 2.34 2.27 239.5 285.0 294.5
NetVLAD [2] 19.5 27.4 32.0 56.5 64.2 63.6 2.85 2.36 2.38 234.0 286.5 294.0
DINO V2 [51] 18.2 26.5 32.1 61.1 65.6 64.8 2.32 2.29 2.31 219.0 270.0 282.5
$R^{2}$ormer [90] 15.2 16.9 26.5 57.9 65.4 65.7 3.03 2.24 2.21 108.0 185.5 214.0
P-NetVLAD [28] 18.9 22.5 29.7 59.7 64.9 64.5 2.67 2.22 2.27 212.5 264.0 283.5
VOP 17.3 22.7 29.2 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
VOP-recall 22.4 31.1 37.6 59.0 64.0 64.1 2.56 2.42 2.49 222.9 286.4 296.2
Table 1: Relative pose estimation on the MegaDepth dataset on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. Standard VOP is trained to maximize the AUC score of the relative pose errors. Version VOP-recall maximizes the recall during training.
@@
AnyLoc [33] 88.9 90.2 90.2 0.76 0.78 0.77 373.2 425.1 438.0
CosPlace [5] 88.3 89.7 89.7 0.77 0.77 0.78 349.7 426.8 437.2
NetVLAD [2] 89.5 90.3 89.9 0.74 0.74 0.79 365.8 147.3 440.1
DINO V2 [51] 87.1 89.2 88.9 0.69 0.67 0.67 338.9 378.2 399.8
$R^{2}$ormer[90] 49.1 74.5 85.2 35.2 6.94 0.83 116.5 299.1 379.6
P-NetVLAD [28] 86.7 88.4 90.2 0.81 0.83 0.81 333.3 412.4 466.8
VOP 83.6 91.1 90.4 0.69 0.61 0.72 293.9 414.7 460.0
Table 2: Relative pose estimation on the ETH3D dataset [67] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. The proposed VOP is trained on the MegaDepth dataset.
@@ 4.2 Generalization Experiments
Moreover, we test the proposed method on multiple real-world scenes from the ETH3D [67] and PhotoTourism [71] datasets for relative pose estimation, and InLoc [74] for localization. We use the AUC-validated model from MegaDepth, i.e., no fine-tuning or retraining were done on the following datasets.
ETH3D. The ETH3D [67] benchmark consists of images with various viewpoints from indoor and outdoor scenes, with ground-truth poses reconstructed by laser point clouds. 13 scenes are used for testing, tens of images from each. All the images in the dataset are used as queries to select database images from the rest. Table 2 shows that the proposed VOP method is always among the top-performing methods, being either the best or second best across all configurations. The overall best results in each metric are always obtained by the proposed VOP. This highlights that VOP performs accurately even without retraining and, thus, is applicable for unseen scenes. Unfortunately, R2Former does not generalize well on ETH3D.
PhotoTourism. Table 3 reports the pose estimation results on 11 testing scenes from PhotoTurism [71], with 100 images in each. We outperform the SOTA competitors in terms of pose estimation accuracy, which shows that the overlap-retrieved images are beneficial in such a geometric pipeline estimating relative poses. Moreover, all Tables 123 show that the number of inliers is not an indicator of accuracy, as we result in fewer inliers while being more accurate than baselines. We have an exceptionally large performance improvement on the top-1 retrieval scores and on the top-5 score on the PhotoTourism dataset.
@@
CosPlace [5] 53.4 60.8 61.2 3.78 2.91 2.87 291.5 347.0 354.1
NetVLAD [2] 55.0 59.3 60.3 3.75 3.10 2.94 283.9 340.0 348.1
DINO V2 [51] 61.1 63.1 63.0 3.17 2.82 2.75 266.0 338.6 356.3
@@
Table 3: Relative pose estimation on the PhotoTourism dataset [71] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. VOP is trained on the MegaDepth dataset.
@@
InLoc. Additionally, we ran all methods on the InLoc dataset [74] (without fine-tuning) using hloc [61] for localization to demonstrate their generalization capabilities in the case of a large domain gap (indoor/outdoor). To retrieve top-40 database images, we prefilter the top-100 by CLS tokens and rerank them by VOP. Columns ’InLoc-DUC1’ and ’InLoc-DUC2’ in Table 4 show the proposed VOP achieves high accuracy, showcasing its generalization abilities.
MegaDepth ETH3D PhotoTourism InLoc-DUC1 InLoc-DUC2 AVG
Metrics AUC@10 on top-5 $\uparrow$ recall@5 on top-40 $\uparrow$
AnyLoc [33] 65.3 90.2 60.3 74.7 75.6 73.2
Cosplace [5] 53.5 89.7 60.8 69.2 74.8 69.6
Cosplacee* [5] 65.8 79.7 65.9 41.4 29.0 56.4
NetVLAD [2] 64.2 90.3 59.3 65.7 71.0 70.1
DINO V2 [51] 65.6 89.2 63.1 63.6 71.0 70.5
VOP 67.6 91.1 67.5 72.2 77.1 75.1
Table 4: Relative pose estimation (AUC@10) on three benchmarks and indoor localization (recall@5) on InLoc. Cosplace* is fine-tuned on Megadepth dataset.
Fairness of Comparison. To demonstrate that the improvements from the proposed VOP do not stem from using different training data, we fine-tuned Cosplace [5] on MegaDepth. As shown in the fourth row (Cosplace*) of Table 4, fine-tuning improves the performance of Cosplace on relative pose estimation on Megadepth and PhotoTourism. However, it largely reduces its accuracy on the ETH3D and InLoc datasets.
4.3 Discussion on Patch-level Descriptors
DINO V2 Features. We investigated the number of patches to be used to split the images. The experiments were conducted on the testing scenes of MegaDepth, with all the images resized to $224$. From the DINO V2 (ViT-14) pretrained model with a patch size of $14^{2}$, we can extract 256 patch descriptors. Then, average pooling is applied to aggregate the patch descriptors to different numbers of patches, such as 4, 16, 64. When the patch number is set to 1, DINO V2 is applied directly to obtain a global image descriptor. Fig. 5 shows the global descriptors (#patches = 1) do not always perform well on the image retrieval problem, as indicated by having a lower recall than when 64 patches are used. This clearly demonstrates that patch-level features have the potential to improve visual localization. Similar to the experiments in the previous sections, the patch-level descriptors learned for our method enhance the original DINO V2 features.
Figure 5: Image retrieval evaluation by DINO V2 backbone with different numbers of patches, e.g., 256. The recalls, AUC@10 scores, and median pose errors in degrees are shown at top-1, 5, and 10 retrievals. When only one patch is used (orange bar), directly DINO V2 features are employed.
Validation Prefilter Augment Dropout AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
VOP Recall $\times$ $\checkmark$ $\times$ 57.4 66.0 63.7 2.87 2.17 2.34 232.5 262 285.5
$\checkmark$ $\checkmark$ $\times$ 57.5 65.2 64.5 2.96 2.19 2.31 235.5 272.5 288.5
AUC $\checkmark$ $\times$ $\times$ 61.9 65.1 64.7 2.37 2.19 2.30 209.0 272.5 282.0
$\times$ $\checkmark$ $\times$ 63.7 66.3 65.3 2.23 2.18 2.31 179.5 222.0 233.0
$\checkmark$ $\checkmark$ $\times$ 64.0 66.7 66.9 2.26 2.18 2.22 213.0 263.0 256.5
$\checkmark$ $\checkmark$ $\checkmark$ 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
Table 5: Relative pose estimation on the MegaDepth dataset [71] on the top-$k$ retrieved images using different configurations, with the best results in bold at each $k$ and the overall best underlined. The validation column refers to the VOP model used. Prefilter indicates if the CLS token was employed to shortlist the potential candidates before overlap prediction. Augment refers to whether data augmentation was used.
Ablations. Table 5 shows the relative pose estimation performance on the top-$k$ images retrieved by the model chosen from the best recall/AUC, with or w/o global prefilter (CLS tokens), and dropout layer. Testing sets of MegaDepth are used, as in Table 1. The CLS prefiltering improves the AUC scores. However, it increases the median pose errors marginally at the same time. As shown in the fifth row, data augmentation is essential in robustly learning the embeddings. Also, the last two rows in Table 4 and Table 5 show that the dropout layer improves the performance on Megadepth and helps with generalization for pose estimation on other data and indoor localization.

### v2

@@ 3.3 Image Retrieval by Overlap Scores
The retrieval pipeline is visualized in Fig. 4. In the online phase, we obtain the patch and, optionally, the global embeddings of the query image. Next, for each query image patch, we perform a radius search in the database of patch embeddings to determine all potentially overlapping candidates. Setting a radius threshold $\epsilon$ is easy as the employed cosine similarity falls between 0 and 1. We tune this threshold on the validation set and use a fixed one in all our experiments. Note that k-nearest-neighbors algorithm is insufficient in our case as we would like to find all potentially overlapping patches to determine image overlaps. Such a radius search can be efficiently performed by standard bounding structures, e.g., kd-tree, and other implementation tricks. Optionally, we also use the global descriptor to prefilter the images to a shortlist.
To obtain the top-$k$ images, we perform voting such that the overlap calculation becomes:
$\hat{O}(I_{i},I_{j})=\sum_{(p,q)\in\mathcal{C}}\llbracket sim(p,q)>\epsilon\rrbracket,$ (7)
@@
$\hat{O}(I_{i},I_{j})=\sum_{(p,q)\in\mathcal{C}}\max(\delta(e_{p},e_{q})-\epsilon,0).$ (8)
Weighted Voting. To highlight the distinguishable query patches, we propose to weight the query patches from each image using a bag of visual words approach [11]. Specifically, we apply TF-IDF [41] weights to all the query patches, assigning lower weights to patches that frequently appear in queries but are rare in the matched database patches. As illustrated below, it is preferable to have more neighbor patches if fewer images share the same neighborhood with the query patch. For each query patch, the weight $t_{i}$ is calculated as follows:
@@
where $n_{id}$ is the number of neighbor patches in database images, defined as the cosine similarities of the patch descriptors within a specified radius, and $n_{d}$ is the total number of patches in each image, i.e., 256 in our case, with a patch size of (14, 14) on the resized image size of 224. Parameter $N$ represents the total number of database images, and $n_{i}$ denotes the number of database images that contain at least one neighbor of the query patch. Patches that frequently appear are deemed less important and thus are down-weighted. Weight $t_{i}$ for each query patch is incorporated into the overlap score computation as follows: $t_{i}*\hat{O}(I_{i},I_{j})$. Note that the map weights are precomputed during the mapping phase.
@@ 4 Experimental Results
Our experiments include estimating two-view epipolar geometry and localizing query images from retrieved images. The VOP model is trained on 153 scenes from the MegaDepth dataset [28], which includes Internet-sourced images and depths reconstructed via Structure-from-Motion (SfM) and multi-view stereo techniques. Beyond the MegaDepth testing scenes, we evaluate the generalization ability of the proposed VOP on 11 scenes from PhotoTourism [49], 13 scenes from ETH3D [45], and the InLoc dataset [51]. We benchmark the proposed method against several competitors, including the DINOv2, CNN-based retrieval methods like NetVLAD [2] and CosPlace [6], as well as the state-of-the-art (SOTA) retrieval method AnyLoc [25] and DINOv2-SALAD [24] utilizing the same DINOv2 backbone as VOP. Besides those global-only methods, we compare VOP with the SOTA reranking-based methods: the ViT-based $R^{2}$ormer [62] and the CNN-based Patch-NetVLAD [20]. Also, we finetune Cosplace* on the same dataset our method was trained on to demonstrate that the improved performance is not due to the different training data. In addition, we test reranking methods [20] [62] on the shortlists created by DINOv2 [CLS] token, similarly as done for VOP, marked with $\dagger$ in the tables.
@@
Training Details. The models underwent training on the 153 MegaDepth scenes, employing contrastive loss to differentiate negative and positive patch pairs. Each scene contributes 150 positive and 150 negative image pairs to the training set, standardized to an image resolution of $224\times 224$ pixels. We use the DINOv2 pre-trained model with a patch size of $14$ to extract $1024$-dimensional descriptors and reduce them to 256 dimensions by a fully-connected layer, i.e., the first layer or our training encoder head. Negative images, characterized by zero overlaps (e.g., sourced from distinct scenes), contrast with positive samples randomly selected from the pairs with 10% to 70% overlapping patches. Notably, the contrastive loss framework is predicated on patch-level negative/positive distinctions rather than entire images. Note that cosine similarity is used in the contrastive loss. Following LightGlue [29], patch-level labels are generated from depth information and visual features, utilizing the GlueFactory library [36].
@@
The training was conducted over 30 epochs with a batch size of 64, employing the validation losses as the primary criterion for best checkpoint selection and utilizing a learning rate of $41e-$ 4. The voting scheme and the radius for neighbor searching were optimized on the validation set by AUC scores, which included a single scene for this purpose. During inference, scenes “0015” (comprising approximately 0.4K images) and “0022” (encompassing approximately 0.5K images) from MegaDepth were designated for testing, enabling a focused evaluation of the model performance on these specific subsets.
Evaluation Metrics. In the evaluation, we assess the relative poses between the query and the top-$k$ retrieved images using robust estimation techniques. This involves deploying SuperPoint [13] and the state-of-the-art LightGlue [29, 44] for feature detection and matching, followed by OpenCV-RANSAC [17] running for 10K iterations to compute the relative poses. The performance is measured against ground truth (GT) poses by calculating the Area Under the Recall curve (AUC) thresholded at $10^{\circ}$ [59, 8], along with the median pose error. The RANSAC threshold is optimized from a set of pixel values $\{0.5,1,2,3\}$ on the training set. Also, we report the average number of inliers across the retrieved image pairs found by RANSAC. Note that the number of inliers is not an indicator of the accuracy of the estimated pose, as discussed and verified in MQNet [4] and FSNet [5]. In scenarios where $k>1$, for each query, we follow the practical approach and select only the image pair that has the highest number of inliers. This comprehensive evaluation allows us to thoroughly assess the performance not just in retrieving relevant images but also in accurately estimating their relative poses.
@@
Additionally, we visualize some retrieval examples with their baselines and accuracy in Fig. 8 to show the aim of designing retrieval methods tailored for geometric challenges, e.g., pose estimation.
@@ 4.1 Pose Estimation on MegaDepth
Table 1 presents the results on the MegaDepth, highlighting the performance of the proposed VOP. The proposed method secures comparable scores for the AUC@10 criterion and exhibits the lowest median pose errors among the SOTA baselines. The underlined results mark the optimal configuration for relative pose accuracy, evidencing the superior performance of the proposed VOP approach when utilizing the top-5 images. This configuration outperforms the global-only retrieval baselines (the first seven rows), reinforcing the significance of selecting correct metrics and model configurations tailored to the specific demands of real-world downstream applications. Our method outperforms the recent reranking approaches $R^{2}$ormer [62] and Patch-NetVLAD [20] on MegaDepth when using their pretrained models. When using the same shortlist we use, their performance (with $\dagger$) is marginally better than ours. The fine-tuned CosPlace (CosPlace*) achieves better accuracy than the pretrained model, especially achieving good results on top-10 pairs. Next, we will show the generalization and robustness of the compared methods on unseen data.
@@
NetVLAD [2] 56.5 64.2 63.6 2.85 2.36 2.38 234.0 286.5 294.0
CosPlace [6] 54.5 53.5 64.3 3.14 2.34 2.27 239.5 285.0 294.5
CosPlace* [6] 59.3 65.8 66.2 2.58 2.65 2.10 164.0 162.0 255.5
DINOv2 [35] 61.1 65.6 64.8 2.32 2.29 2.31 219.0 270.0 282.5
AnyLoc [25] 60.8 65.3 64.1 2.38 2.26 2.35 220.5 273.0 287.5
SALAD [24] 53.9 61.8 62.6 3.23 2.61 2.54 224.5 278.0 285.0
P-NetVLAD [20] 59.7 64.9 64.5 2.67 2.22 2.27 212.5 264.0 283.5
[…]
$R^{2}$ormer [62] 57.9 65.4 65.7 3.03 2.24 2.21 108.0 185.5 214.0
[…]
VOP 61.8 67.6 66.0 2.36 2.02 2.28 177.5 246.5 251.0
Table 1: Relative pose estimation on MegaDepth using top-$k$ images retrieved by different methods, with the best results in bold at each $k$ and the overall best underlined. results are tested using reranking methods prefiltered by the [CLS] tokens from DINOv2.
@@ 4.2 Generalization Experiments
We test the proposed method on multiple real-world scenes from the ETH3D [45] and PhotoTourism [49] datasets for relative pose estimation, and InLoc [51] for localization. We use the model trained on MegaDepth, i.e., no fine-tuning or retraining is done on the following datasets.
@@
NetVLAD [2] 89.5 90.3 89.9 0.74 0.74 0.79 365.8 147.3 440.1
CosPlace [6] 88.3 89.7 89.7 0.77 0.77 0.78 349.7 426.8 437.2
CosPlace* [6] 43.6 79.7 85.9 38.3 0.79 0.76 162.6 279.0 360.5
DINOv2 [35] 87.1 89.2 88.9 0.69 0.67 0.67 338.9 378.2 399.8
AnyLoc [25] 88.9 90.2 90.2 0.76 0.78 0.77 373.2 425.1 438.0
SALAD [24] 89.6 90.2 89.9 0.73 0.76 0.78 385.9 411.2 442.0
P-NetVLAD [20] 86.7 88.4 90.2 0.81 0.83 0.81 333.3 412.4 466.8
P-NetVLAD [20] 80.5 87.7 89.8 0.87 0.65 0.66 273.2 319.1 355.3
$R^{2}$ormer[62] 49.1 79.8 90.1 35.2 6.94 0.83 116.5 252.8 364.9
$R^{2}$ormer[62] 50.6 81.5 88.2 27.3 0.87 0.80 137.0 318.3 401.5
VOP 83.6 91.1 90.4 0.69 0.61 0.72 293.9 414.7 460.0
Table 2: Relative pose estimation on the ETH3D dataset [45] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. The proposed VOP is trained on the MegaDepth dataset. results are tested using reranking methods prefiltered by the [CLS] tokens.
ETH3D [45] benchmark consists of images with various viewpoints from indoor and outdoor scenes, with ground-truth poses reconstructed by laser scanner. 13 scenes are used for testing. All images in the dataset are used as queries to select database images from the rest. Table 2 shows that VOP is always among the top-performing methods. The overall best results in all accuracy metrics are always obtained by VOP by selecting top-5 images. This highlights that VOP performs accurately even without retraining and, thus, is applicable to unseen scenes. The fine-tuned Cosplace does not generalize well to ETH3D, and yields degraded results, especially on the top-1. The reranking methods, P-NetVLAD and $R^{2}$ormer, significantly lag behind the performance of VOP, both in their original version and when combined with DINOv2 [CLS] tokens.
PhotoTourism. Table 3 reports the relative pose estimation results on 11 testing scenes from the PhotoTurism [49] dataset. The proposed VOP achieves the lowest median pose error and the second-best AUC scores. Similar to Tables 1 and 2, Table 3 shows that the number of inliers is not an indicator of accuracy, as we result in fewer inliers while being more accurate than baselines.
AUC@10 $\uparrow$ Med. pose error ($\downarrow$ # inliers $\uparrow$
Top $\rightarrow$ 1 5 10 1 5 10 1 5 10
NetVLAD [2] 55.0 59.3 60.3 3.75 3.10 2.94 283.9 340.0 348.1
CosPlace [6] 53.4 60.8 61.2 3.78 2.91 2.87 291.5 347.0 354.1
CosPlace* [6] 58.3 65.9 66.1 3.59 2.51 2.43 145.9 247.9 281.8
DINOv2 [35] 61.1 63.1 63.0 3.17 2.82 2.75 266.0 338.6 356.3
@@
SALAD [24] 53.5 60.1 59.6 4.07 3.19 3.17 270.7 338.1 349.7
P-NetVLAD [20] 61.9 62.4 61.8 2.76 2.65 2.76 261.9 315.2 339.4
P-NetVLAD [20] 62.1 63.6 62.8 2.69 2.63 2.71 246.5 322.8 338.5
@@
Table 3: Relative pose estimation on the PhotoTourism dataset [49] on the top-$k$ retrieved images using different retrieval methods, with the best results in bold at each $k$ and the overall best underlined. VOP is trained on the MegaDepth dataset. * shows fine-tuned Cosplace on Megadepth.
Methods
NetVLAD
 [2]
Cosplace
 [6]
Cosplace*
 [6]
DINOv2
 [35]
AnyLoc
 [25]
SALAD
 [24]
[…]
 [20]
[…]
 [62]
VOP
DUC1 65.7 69.2 41.4 63.6 74.7 71.2 60.1 47.0 72.2
DUC2 71.0 74.8 29.0 71.0 75.6 78.6 55.0 66.4 77.1
[…]
Methods
NetVLAD
 [2]
Cosplace
 [6]
Cosplace*
 [6]
DINOv2
 [35]
AnyLoc
 [25]
SALAD
 [24]
[…]
 [20]
$\dagger R^{2}$ormer
 [62]
VOP
Avg. Accuracy (%) $\uparrow$ 70.1 69.6 56.4 70.5 73.2 72.3 66.2 65.3 75.1
Avg. med. pose error ($\downarrow$ 12.1 12.0 12.0 11.9 12.0 12.2 11.9 11.8 11.5
Table 5: Average accuracy (in percentage) and median pose error (in degrees) on the retrieved top-5 images on the Megadepth [28], ETH3D [45], PhotoTourism [49], and InLoc [51] datasets. InLoc is excluded from the median error, as the official website does not report it. indicates prefiltering by the [CLS] tokens. * shows fine-tuned Cosplace on Megadepth. The best results are shown in bold.
@@ Methods
InLoc. Additionally, we ran all methods on the InLoc dataset [51] (without fine-tuning) using hloc [44] for localization to show their generalization capabilities in the case of a large domain gap (in/outdoor). Following hloc [44], we retrieve top-40 database images for all the retrieval methods. For the reranking ones (including VOP), we rerank the top-100 shortlist created by [CLS] token. Table 4 shows the proposed VOP achieves the second highest accuracy, showcasing its generalization abilities. Note that we only show results for P-NetVLAD and $R^{2}$ormer with DINOv2 [CLS] token as their original version fails on this dataset.
In summary, we show the average accuracy and median pose errors averaged over all tested datasets in Table 5. Note that InLoc is excluded from the median pose error, as the official website does not report it. Overall, the proposed VOP leads to the highest accuracy and lowest pose errors. Fine-tuning Cosplace on Megadepth reduces its accuracy on average, demonstrating that our improved performance is not simply due to the different training data. Again, we do not report results for P-NetVLAD and $R^{2}$ormer without DINOv2 [CLS] token due to their failure on InLoc.
Figure 6: AUC@10 scores on the queries with top-5 images from database images lower than the limits along the horizontal axis.
4.3 Discussions
Difficult Scenario Experiment. We conducted an experiment on MegaDepth to retrieve the top-5 images in the low-overlap scenario. For this test, for each query image, we prefiltered the database by keeping only those images that have a lower overlap with the query than a threshold. Then, we ran image retrieval on this prefiltered database. As shown in Fig. 6, VOP demonstrates clear improvements compared with other methods in this challenging scenario.